# latex‘h’ float specifier changed to ‘ht’

### v1

@@ Abstract
Sequential recommendation aims to estimate dynamic user preferences and sequential dependencies among historical user behaviors. Attention-based models have proven effective for sequential recommendation, but they suffer from inference inefficiency due to the quadratic computational complexity of attention mechanisms, particularly for long-range behavior sequences. Inspired by the recent success of state space models (SSMs) in control theory, which provide a robust framework for modeling and controlling dynamic systems, we present EchoMamba4Rec. Control theory emphasizes the use of SSMs for managing long-range dependencies and maintaining inferential efficiency through structured state matrices. EchoMamba4Rec leverages these control relationships in sequential recommendation and integrates bi-directional processing with frequency-domain filtering to capture complex patterns and dependencies in user interaction data more effectively. Our model benefits from the ability of state space models (SSMs) to learn and perform parallel computations, significantly enhancing computational efficiency and scalability. It features a bi-directional Mamba module that incorporates both forward and reverse Mamba components, leveraging information from both past and future interactions.
Additionally, a filter layer operates in the frequency domain using learnable Fast Fourier Transform (FFT) and learnable filters, followed by an inverse FFT to refine item embeddings and reduce noise. We also integrate Gate Linear Units (GLU) to dynamically control information flow, enhancing the model’s expressiveness and training stability. Experimental results demonstrate that EchoMamba significantly outperforms existing models, providing more accurate and personalized recommendations. This study underscores the potential of integrating bi-directional and frequency-domain techniques with control theory to advance the state-of-the-art in sequential recommendation systems. Our code can be found at https://github.com/wyd0042/EchoMamba4Rec.
@@ 3.5 Filter Layer
The second reason we named our model EchoMamba4Rec is due to our use of a Fourier Transform-based (FFT) filter [30] to capture sequential dependencies. In the filter layer, we perform a filtering operation for each feature dimension in the frequency domain, followed by skip connections and layer normalization. Given the input item representation matrix $\mathbf{F}^{l}\in\mathbb{R}^{n\times d}$ at layer $l$ (where $l=0$ implies $\mathbf{F}^{0}=\mathbf{E_{I}}$), where $n$ is the number of items and $d$ is the feature dimension, we first perform a one-dimensional FFT along the item dimension to convert $\mathbf{F}^{l}$ into the frequency domain:
@@
where $\mathcal{F}(\cdot)$ denotes the one-dimensional FFT, and $\mathbf{X}^{l}$ is a complex tensor representing the spectrum of $\mathbf{F}^{l}$. We then modulate this spectrum by multiplying it with a learnable filter $\mathbf{K}\in\mathbb{C}^{n\times d}$:
@@ 4.1 Experimental Setup
Datasets. Two datasets contain product reviews and ratings for the “Beauty” and “Video Games” categories on Amazon (Amazon-Beauty and Amazon-Video-Game)[31], which analyzed using the RecBole framework [32]. For each user, the interaction sequence is constructed by sorting their interaction records according to their timestamps. Users and items with less than 5 interactions are filtered. The preprocessed statistics of each dataset are summarized in Table 1.
Table 1: Statistics of the Amazon Review Datasets for Beauty and Video Games categories. These datasets provide comprehensive user-item interaction data, facilitating research on sequential recommendation systems. The statistics include the number of users, items, interactions, and the average interaction length per user.
@@
Beauty 22,363 12,101 198,502 8.9
Video-Games 14,494 6,950 132,209 9.1
@@
Method Amazon-Beauty Amazon-Video-Games
HR@10 NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10
NARM 0.0627 0.0347 0.0262 0.1032 0.0530 0.0379
GRU4Rec 0.0606 0.0332 0.0249 0.1030 0.0536 0.0380
SASRec 0.0739 0.0404 0.0302 0.1203 0.0561 0.0366
BERT4Rec 0.0711 0.0426 0.0309 0.1106 0.0508 0.0324
Mamba4Rec 0.0621 0.0369 0.0292 0.1057 0.0587 0.0441
Bi-Mamba4Rec 0.0781 0.0423 0.0313 0.1200 0.0580 0.0392
Bi-Mambaformer4Rec 0.0773 0.0304 0.0412 0.1205 0.0451 0.0626
EchoMamba4Rec 0.0833 0.0438 0.0318 0.1254 0.0633 0.0445
4.2 Recommendation performance
Based on the results presented in Table 2, the evaluation of models reveals notable trends. RNN-based models, including NARM and GRU4Rec, demonstrate competitive but relatively lower performance compared to attention-based models such as SASRec and BERT4Rec. And the Mamba model variants, outperform both RNN and attention-based models across both Amazon-Beauty and Amazon-Video-Games datasets. Notably, EchoMamba4Rec achieves superior performance compared to baselines based on RNN, Transformer, Mamba and thier hybrid showcasing its effectiveness in capturing intricate user-item interactions and providing accurate recommendations.

### v2

@@ Abstract
Predicting user preferences and sequential dependencies based on historical behavior is the core goal of sequential recommendation. Although attention-based models have shown effectiveness in this field, they often struggle with inference inefficiency due to the quadratic computational complexity inherent in attention mechanisms, especially with long-range behavior sequences. Drawing inspiration from the recent advancements of state space models (SSMs) in control theory, which provide a robust framework for modeling and controlling dynamic systems, we introduce EchoMamba4Rec. Control theory emphasizes the use of SSMs for managing long-range dependencies and maintaining inferential efficiency through structured state matrices. EchoMamba4Rec leverages these control relationships in sequential recommendation and integrates bi-directional processing with frequency-domain filtering to capture complex patterns and dependencies in user interaction data more effectively. Our model benefits from the ability of state space models (SSMs) to learn and perform parallel computations, significantly enhancing computational efficiency and scalability. It features a bi-directional Mamba module that incorporates both forward and reverse Mamba components, leveraging information from both past and future interactions.
Additionally, a filter layer operates in the frequency domain using Fast Fourier Transform (FFT) and learnable filters, followed by an inverse FFT to refine item embeddings and reduce noise. We also integrate Gate Linear Units (GLU) to dynamically control information flow, enhancing the model’s expressiveness and training stability. Experimental results demonstrate that EchoMamba significantly outperforms existing models, providing more accurate and personalized recommendations. This study underscores the potential of integrating bi-directional and frequency-domain techniques with control theory to advance the state-of-the-art in sequential recommendation systems. Our code can be found at https://github.com/wyd0042/EchoMamba4Rec.
@@ 3.5 Filter Layer
The second reason we named our model EchoMamba4Rec is due to our use of a Fourier Transform-based (FFT) filter [30] to capture sequential dependencies. In the filter layer, a filtering operation is performed for each feature dimension in the frequency domain, followed by the application of skip connections and layer normalization. Given the input item representation matrix $\mathbf{F}^{l}\in\mathbb{R}^{n\times d}$ at layer $l$ (where $l=0$ implies $\mathbf{F}^{0}=\mathbf{E_{I}}$), where $n$ is the number of items and $d$ is the feature dimension, we first perform a one-dimensional FFT along the item dimension to convert $\mathbf{F}^{l}$ into the frequency domain:
@@
where $\mathcal{F}(\cdot)$ denotes the one-dimensional FFT, and $\mathbf{X}^{l}$ is a complex tensor representing the spectrum of $\mathbf{F}^{l}$. We then modulate this spectrum by multiplying it with a filter $\mathbf{K}\in\mathbb{C}^{n\times d}$ which is learnable:
@@ 4.1 Experimental Setup
Datasets. Three datasets contain product reviews and ratings for the “Beauty”, “Video Games” categories on Amazon (Amazon-Beauty and Amazon-Video-Games) [31] and movie ratings from MovieLens (MovieLens-1M) [32], which analyzed using the RecBole framework [33]. For each user, the interaction sequence is constructed by sorting their interaction records according to their timestamps. Users and items with less than 5 interactions are filtered. The preprocessed statistics of each dataset are summarized in Table 1.
Table 1: Statistics of the Amazon Review Datasets for Beauty and Video Games categories and MovieLens Datasets for movie ratings. These datasets provide comprehensive user-item interaction data, facilitating research on sequential recommendation systems. The statistics include the number of users, items, interactions, and the average interaction length per user.
@@
MovieLens-1M 6,040 3,416 999,611 165.4
Amazon-Beauty 22,363 12,101 198,502 8.9
Amazon-Video-Games 14,494 6,950 132,209 9.1
@@
4.2 Recommendation performance
Based on the results presented in Table 2, the evaluation of models reveals notable trends. RNN-based models, including NARM and GRU4Rec, demonstrate competitive but relatively lower performance compared to attention-based models such as SASRec and BERT4Rec. And the Mamba model variants, outperform both RNN and attention-based models across MovieLens-1M, Amazon-Beauty, and Amazon-Video-Games datasets. Notably, EchoMamba4Rec achieves superior performance compared to baselines based on RNN, Transformer, Mamba and thier hybrid showcasing its effectiveness in capturing intricate user-item interactions and providing accurate recommendations.
Method MovieLens-1M Amazon-Beauty Amazon-Video-Games
HR@10 NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10
NARM 0.2884 0.1445 0.1267 0.0640 0.0369 0.0287 0.1057 0.0556 0.0405
GRU4Rec 0.2642 0.1464 0.1106 0.0664 0.0365 0.0274 0.1101 0.0573 0.0414
SASRec 0.2429 0.1262 0.0905 0.0739 0.0404 0.0302 0.1203 0.0561 0.0366
BERT4Rec 0.2910 0.0979 0.0663 0.0711 0.0426 0.0309 0.1106 0.0508 0.0324
Mamba4Rec 0.3069 0.1509 0.1223 0.0621 0.0369 0.0292 0.1057 0.0587 0.0441
Bi-Mamba4Rec 0.3051 0.1485 0.1398 0.0781 0.0423 0.0313 0.1200 0.0580 0.0392
Bi-Mambaformer4Rec 0.3166 0.1470 0.1106 0.0773 0.0304 0.0412 0.1205 0.0451 0.0626
EchoMamba4Rec 0.3215 0.1529 0.1458 0.0833 0.0438 0.0318 0.1254 0.0633 0.0445